OpenStudy (anonymous):

Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 6-inch apothem A = sq. in.

OpenStudy (sriram):

side of triangle=2A/tan(30) =12*sqrt3 area=(side)^2*cos30 =144*3*(sqrt3)/2